OpenStudy (anonymous):

Can someone help me add/subtract and simplify the radicals? I'll comment the question I'm stuck on, and can someone show me how to do it, rather than just give me the answer?

OpenStudy (anonymous):

\[\sqrt{6}+\sqrt{54}+2\sqrt{24}\]

OpenStudy (owlcoffee):

Hello!, I'll give you some theory and tricks to simplify these radicals.

OpenStudy (anonymous):